# ReUseIt: Synthesizing Reusable AI Agent Workflows for Web Automation

### v1

@@ Abstract.
AI-powered web agents have the potential to automate repetitive tasks, such as form filling, information retrieval, and scheduling, but they struggle to reliably execute these tasks without human intervention, requiring users to provide detailed guidance during every run. We address this limitation by automatically synthesizing reusable workflows from an agent’s successful and failed attempts. These workflows incorporate execution guards that help agents detect and fix errors while keeping users informed of progress and issues. Our approach enables agents to successfully complete repetitive tasks of the same type with minimal intervention, increasing the success rates from 24.2% to 70.1% across fifteen tasks. To evaluate this approach, we invited nine users and found that our agent helped them complete web tasks with a higher success rate and less guidance compared to two baseline methods, as well as allowed users to easily monitor agent behavior and understand failures.
@@ 1. Introduction
Data scientists, business analysts, and everyday users often need to perform repetitive interactions with the web that they may wish to automate. For example, a data scientist who wants to analyze housing markets needs to periodically scrape data from multiple online housing listings; a market researcher who hopes to understand changing customer attitudes towards certain products needs to collect information from multiple online forums; and, a customer who wishes to optimize their expenses to purchase a popular item needs to visit different websites to check availability and cost. AI-powered web agents have the potential to assist with these user needs that entail repetitive web tasks. Powered by vision language models (VLMs), web agents can perceive and take human-like actions on the web (OpenAI, 2025b; Mozannar et al., 2025d). With high-level inputs (prompts) from the user, web agents can autonomously interact with the web to navigate pages, submit forms, and collect information, even when they encounter dynamic web content (e.g., websites with changing layouts, pop-up menus) that traditional script-based automation techniques cannot handle (Dong et al., 2022).
@@
Figure 2. Given a user instruction, the web agent can make low-level but critical mistakes that prevent it from completing tasks. For example, the agents mistakenly used the ‘+’ next to children to add adults on Booking.com (left) and selected ‘round-trip’ from the drop-down menu despite users asking for ‘one-way’ flights on Google Flights (right).
@@
However, despite their impressive capability, existing web agents struggle to reliably complete web tasks and apply a workflow for future reuse. From a high-level instruction, prior evaluation has shown that the state-of-the-art web agents achieved 66.0% (Anthropic Sonnet 3.7 CUA) and 59.8% (OpenAI CUA) accuracy to execute common web tasks (Singh, 2025). Even if the agent is provided with a successful workflow (e.g., guided by users during execution (Mozannar et al., 2025d)), the agent struggles to reapply the workflow to the same task at a different time or to new tasks with similar workflows. In our benchmark evaluation of six common tasks, the average success rate of repeating the same tasks (five times each task) was 28.0% and generalization to related tasks (five times each task) was 22.5%, using the Magentic-UI agent framework (Mozannar et al., 2025d). The agent would sometimes hang, repetitively attempt wrong actions, or misunderstand the status of the workflow. Thus, to use agents for web automation, users currently need to interactively correct the agent’s workflow every time they execute the web task, which presents a substantial limitation for the desired goal of automation. In fact, it can be difficult for users to watch agent actions on screen to effectively understand the progress of the AI agent towards their given goals and provide feedback when they get stuck. Coupled with the fact (based on our formative user study, Section 3.2) that users often struggle at pinpointing and offering guidance to resolve low-level agent mistakes (e.g., cannot locate the right button to update the number of adult passengers, despite already understanding that it needs to update the passenger number, as illustrated in Figure 2), there are still substantial challenges to using AI agents for web task automation.
@@
To address this, we propose an automatic workflow synthesis approach, ReUseIt , that can produce reliably reusable workflows for AI agents. Motivated by prior research on learning and reusing workflows to improve agent execution performance (Wang et al., 2024; Li et al., 2024; Yu et al., 2025), our approach lets web agents automatically compose workflows from their successful and failed attempts with variational inputs that can be both (1) reliably used towards future repetitive tasks and (2) interpretable for users. Given a user input, such as Figure 1 A, instead of attempting to just complete the task, ReUseIt first runs the task multiple times in parallel with different input variations that are generated based on the potential future reuse of the workflow, as shown in Figure 1 B.1 (in our experiments, the agent first repeats nn=5 for the initial inputs and runs another nn=5 for each different input variation). While the agent often cannot successfully complete all exploration runs, both successful and failed attempts provide valuable information that surfaces potential mistakes the agent can make during executions, especially common pitfalls, and recovery strategies (Figure 1 B.2). Based on these executions, ReUseIt synthesizes an annotated workflow with execution guards consisting of condition checks and fallback actions that describe the website state that the agent is expected to achieve before/after key interaction steps (identified based on prior failed execution traces) and what the agent should do if they find themselves in a different browser state (extracted according to successful experiences), as shown in Figure 1 B.3. With the synthesized workflow, users can apply it to reliably execute new tasks on different websites, as demonstrated in Figure 1 C. The execution guards also serve as the interface between the user and the agent, where users can examine the differences between the saved snapshot and the current workflow to track the progress and provide feedback. With this design, the burden on the user to track agent actions through lengthy action replays is turned into verification of key webpage states, providing users with a better experience in interacting with web agents.
@@
To evaluate the effectiveness of ReUseIt to generate reliable and interpretable workflows for web automation, we used Magentic-UI (Mozannar et al., 2025d) as the agent framework and conducted (1) a benchmark evaluation on whether ReUseIt ’s synthesized workflows can be reliably repeated on both the same and (unseen) task variations on fifteen tasks (three variations per task), and (2) a user study with nine participants to examine whether execution guards in the synthesized workflow can help users to understand and steer web agents. From the benchmark evaluation, web agents guided by the ReUseIt ’s synthesized workflows achieved a 70.1% success rate when executing repetitive tasks, improving the performance by 45.9% over executing without the workflows. From the user study, the execution guards effectively helped users to identify and understand agent issues, and boosted their willingness and trust to use web agents in their own tasks. In sum, this work has three main contributions:
@@ 2. Related Work
LLM-based web agents. LLM-based web agents are autonomous systems that leverage LLMs to navigate websites and execute tasks based on natural language instructions. These agents translate the LLM’s text outputs into web actions (e.g., click links and buttons, fill in forms, or web search) by specialized prompting and tool-use frameworks (OpenAI, 2021, 2025b). These agents have been explored in a wide range of use scenarios, aimed at both professionals and end-users. On the professional side, web agents are being tested as digital “office workers” that could automate data science and business workflows. For example, a recent benchmark, TheAgentCompany (Xu et al., 2025), simulated an enterprise intranet environment with internal web apps and data and evaluated LLM-based agents on everyday workplace tasks. On the consumer side, web agents can serve as advanced personal assistants that perform common web tasks on behalf of users. For instance, a web agent can book hotel rooms, search flight tickets, and reserve restaurants (Yang et al., 2024; Zhou et al., 2023). More use cases include tracking product prices on e-commerce websites, retrieving answers from multiple forums, and summarizing daily feed on social media (Uzair, 2023; OpenAI, 2025a).
@@
Despite rapid progress, LLM-based web agents face significant challenges in repetitive tasks. A common concern is the reliability of task performance. Recent studies have found that only about 30% of multi-step web tasks can be completed successfully by advanced agents like Operator (OpenAI, 2025b), and many other agents achieve far lower success rates on average, often below 20% on realistic web automation benchmarks (Ning et al., 2025; Xue et al., [n. d.]; Singh, 2025; Song et al., 2025). Another fundamental issue is inconsistent behavior that limits reusability. Since agent decisions emerge from probabilistic language model outputs, the exact sequence of actions can vary from run to run, and these agents fall short in being reused in similar tasks (Xue et al., [n. d.]; OpenAI, 2025b; Pan et al., 2024; Levy et al., 2024). Such ad-hoc behavior also leads to unpredictable failures. When an unexpected webpage status change happens (e.g., a pop-up menu), the agent behaves differently across runs and fails in various ways (Ledel, 2025). To address these limitations, we focus on augmenting reliability and consistency with execution guards and self-recovery, and boosting reusability through successful and failed experiences from multiple executions of the same and similar tasks.
@@
Workflow synthesis for web agents. A promising approach to address web agents’ limitations is to automatically learn and reuse execution workflows as part of their reasoning process. For example, Agent Workflow Memory (AWM) (Wang et al., 2024) induces commonly used workflows from an agent’s past experiences and injects them into the agent’s memory to guide future actions. Leveraging learned workflows, AWM significantly boosted success rates on long-horizon web tasks (e.g., +24.6% on the Mind2Web benchmark (Deng et al., 2023)) while reducing the number of execution steps. Other work has explored representing workflows in various forms to guide the agent, including natural language (Li et al., 2024), descriptive workflow language (Deng et al., 2023), and pseudo-code (Zhou et al., 2023; Wang et al., 2024). In our work, we similarly induce agent workflows automatically, with a specific focus on synthesizing execution guards in the workflow to enhance agent reliability and reusability using any representation of the workflow.
@@
Programming by demonstration for web automation. Programming by demonstration (PbD) techniques have also been applied to address web automation challenges via scripting. In this paradigm, a human user performs a web task step-by-step, and the system synthesizes a (deterministic) program that replicates those actions. For example, WebRobot (Dong et al., 2022) constructs web automation scripts by observing user interactions with a web browser. DiLogics (Pu et al., 2023) leverages natural language processing to generalize user demonstrations into web automation scripts that handle diverse input conditions. MIWA (Chen et al., 2023) introduces a mixed-initiative interface for web automation, providing visual explanations and interactive debugging tools to help users refine and trust the generated scripts. In particular, programming by navigation synthesis (Lubin et al., 2025) has been proposed to support the iterative user specification refinement process.
@@
User interaction with web agents. Researchers have proposed keeping humans in the loop during agent execution to monitor and steer the agents. A recent system, Magentic-UI (Mozannar et al., 2025d), exemplifies this line of research. It provides a human-centered interface for web-based agentic systems, allowing users to observe agent actions and intervene during execution. In addition, Uncertainty-Aware GUI Agent (Hao et al., 2025) can detect agent uncertainty in decision phases and request user feedback in ambiguous states; the Dark Patterns study (Tang et al., 2025a) highlights how agents can be misled by deceptive interfaces, and shows that human oversight reduces errors albeit with attention costs; the Human-Centered Evaluation framework (Chen et al., 2025) suggests incorporating human evaluators into GUI agent assessment pipelines to ensure trustworthiness and alignment. By combining human oversight with AI-driven automation, these human-in-the-loop systems and frameworks have demonstrated improved user understanding and agent safety.
Following this idea, our work loops users into the agent’s execution process, such that users can identify and understand agent issues as they happen. Specifically, when the agent is uncertain about how to proceed or encounters a failure that it cannot recover from, it invites user guidance with issue analysis, aiming for more interpretable, trustworthy, and controllable agents.
@@ 3. Motivation
3.1. Evaluation: Reliability of web agents to execute common web interaction tasks
To quantify overall agent performance on common web automation tasks, we collected six tasks from prior work (He et al., 2024; Deng et al., 2025; Krosnick and Oney, 2021; Dong et al., 2022). These tasks require the agents to navigate the web, fill forms, filter, extract information, search, and summarize content, and they span different application domains, including flight reservation, housing, pet, online shopping, research publications, and news. More details on the study tasks are in Appendix A.1. We evaluated the agent’s performance to complete given tasks under three conditions, to evaluate both agents’ base ability to complete the task and their ability to reapply previous low-level traces and high-level plans towards new tasks:
  • Task-Only: The agent is only provided with the task prompt from the user (e.g., “help me search for product information”);
  • Task + Success-Traces: The agent is provided with the user prompt describing the task together with a previous successful execution trace containing low-level, detailed actions taken on each webpage (e.g., opened a drop-down menu, typed content in a text field; Figure 3 A);
  • Task + Magentic-UI Plan: The agent is provided with the task prompt together with a high-level workflow outlining the main steps obtained with Magentic-UI’s plan-learning method (e.g., step 1: navigate to a user-specified website, step 2: search content; Figure 3 B).
To measure the agent’s performance, we used the success rate ($SR$) as the evaluation metric. $SR$ is defined as the number of successful executions out of the total number of runs. For each task, we ran five executions and computed the fraction of successful runs. We also generated variation tasks (with variations on task parameter, category, and websites; more details are presented in Section 4.1) and executed each variation five times to test workflow reuse on closely related tasks. In total, we executed each task and its variations 20 times and report the mean $\pm$ std $SR$ (in $\%$) across all these executions. To determine whether the executions were successful, we manually inspected the screen recordings to judge if the agent had satisfied all user requirements and found the intended results.

### v2

@@ Abstract.
AI-powered web agents have the potential to automate repetitive tasks, such as form filling, information retrieval, and scheduling, but they struggle to reliably execute these tasks without human intervention, requiring users to provide detailed guidance during every run. We address this limitation by automatically synthesizing reusable workflows from an agent’s successful and failed attempts. These workflows incorporate execution guards that help agents detect and fix errors while keeping users informed of progress and issues. Our approach enables agents to successfully complete repetitive tasks of the same type with minimal user intervention, increasing the success rates from 24.2% to 70.1% across fifteen tasks. To evaluate this approach, we invited nine users and found that our agent helped them complete web tasks with a higher success rate and less guidance compared to two baseline methods, as well as allowed users to easily monitor agent behavior and understand its failures.
AI agent reusability and interpretability, workflow synthesis
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789083isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Human computer interaction (HCI)
@@
This figure shows a flowchart for ReUseIt. Starting from the left, ”A User Query” specifies a workflow synthesis query from the user. This query is passed to ”B Web Agent Automatically Synthesizes the ReUseIt Workflow”. In the workflow synthesis process, the web agent first ”B.1 Execute Task Variations Multiple Times”, then ”B.2 Obtain Failed and Successful Executions”, and finally obtains ”B.3 ReUseIt Workflow”. In the last step, the synthesized workflow is applied by the user to ”C Tasks with New Parameters on Different Websites”, for reliable web agent task execution.
@@ 1. Introduction
Figure 2. Given user instructions, web agents can make low-level but critical mistakes that prevent them from completing tasks. For example, the agent mistakenly used the “+” next to children to add adults on Booking.com (left), and mistakenly selected “round trip” from the drop-down menu despite the user asking for “one-way” flights on Google Flights (right).
This figure shows two examples of web agent mistakes. On the left, the web agent was asked to ”search for one-way flights from LA to Chicago for three adults on December 10 using Booking.com.” However, the agent mistakenly added one child, even though the task did not involve any child among the passengers. On the right, the web agent was asked to ”search for one-way flights from Miami to LA for two adults on August 5 using Google Flights.” However, the web agent mistakenly selected round-trip, which contradicts the required one-way flights.
Data scientists, business analysts, and everyday users often need to perform repetitive interactions with the web that they may wish to automate. For example, a data scientist who wants to analyze housing markets needs to periodically scrape data from multiple online housing listings; a market researcher who hopes to understand changing customer attitudes towards certain products needs to collect information from multiple online forums; and, a customer who wishes to optimize their expenses to purchase a popular item needs to visit different websites to check availability and cost. AI-powered web agents have the potential to assist with these user needs that entail repetitive web tasks. Powered by vision language models (VLMs), web agents can perceive and take human-like actions on the web (OpenAI, 2025b; Mozannar et al., 2025a). With high-level inputs (prompts) from the user, web agents can autonomously interact with the web to navigate pages, submit forms, and collect information, even when they encounter dynamic web content (e.g., websites with changing layouts, pop-up menus) that traditional script-based automation techniques cannot handle (Dong et al., 2022).
@@
However, despite their impressive capability, existing web agents struggle to reliably complete web tasks and apply a workflow for future reuse. From a high-level instruction, prior evaluation has shown that the state-of-the-art web agents achieved 66.0% (Anthropic Sonnet 3.7 CUA) and 59.8% (OpenAI CUA) accuracy to execute common web tasks (Singh, 2025). Even if the agent is provided with a successful workflow (e.g., guided by users during execution (Mozannar et al., 2025a)), the agent struggles to reapply the workflow to the same task at a different time or to new tasks with similar workflows. In our benchmark evaluation of six common tasks, the average success rate of repeating the same tasks (five times each task) was 28.0% and generalization to related tasks (five times each task) was 22.5%, using the Magentic-UI agent framework (Mozannar et al., 2025a). The agent would sometimes hang, repetitively attempt wrong actions, or misunderstand the status of the workflow. Thus, to use agents for web automation, users currently need to interactively correct the agent’s workflow every time they execute the web task, which presents a substantial limitation for the desired goal of automation. In fact, it can be difficult for users to watch agent actions on screen to effectively understand the progress of AI agents towards their given goals and provide feedback when they get stuck. Coupled with the fact (based on our formative user study, Section 3.2) that users often struggle at pinpointing and offering guidance to resolve low-level but critical agent mistakes (e.g., cannot locate the right button to update the number of adult passengers, despite already understanding that it needs to update the passenger number, as illustrated in Figure 2), there are still substantial challenges to using AI agents for web task automation.
@@
To address these challenges, we propose an automatic workflow synthesis approach, ReUseIt , that can produce reliably reusable workflows for AI agents. Motivated by prior research on learning and reusing workflows to improve agent execution performance (Wang et al., 2024b; Li et al., 2024; Yu et al., 2025), our approach lets web agents automatically compose workflows from their successful and failed attempts with variational inputs that can be both (1) reliably used towards future repetitive tasks and (2) interpretable for users. Given a user input, such as Figure 1 A, instead of attempting to just complete the task, ReUseIt first runs the task multiple times in parallel with different input variations that are generated based on the potential future reuse of the workflow, as shown in Figure 1 B.1 (in our experiments, the agent first repeats nn=5 for the initial inputs and runs another nn=5 for each different input variation). While the agent often cannot successfully complete all exploration runs, both successful and failed attempts provide valuable information that surfaces potential mistakes the agent can make during executions, especially common pitfalls, and recovery strategies (Figure 1 B.2). Based on these executions, ReUseIt synthesizes an annotated workflow with execution guards consisting of condition checks and fallback actions that describe the website state that the agent is expected to achieve before/after key interaction steps (identified based on prior failed execution traces) and what the agent should do if they find themselves in a different browser state (extracted according to successful experiences), as shown in Figure 1 B.3. With the synthesized workflow, users can apply it to reliably execute new tasks on different websites, as demonstrated in Figure 1 C. The execution guards also serve as the interface between the user and the agent, where users can examine the differences between the saved snapshot and the current workflow to track the progress and provide feedback. With this design, the burden on the user to track agent actions through lengthy action replays is turned into verification of key webpage states, providing users with a better experience in interacting with web agents.
@@
To evaluate the effectiveness of ReUseIt to generate reliable and interpretable workflows for web automation, we used Magentic-UI (Mozannar et al., 2025a) as the agent framework and conducted (1) a benchmark evaluation on whether ReUseIt ’s synthesized workflows can be reliably repeated on both the same and (unseen) task variations on fifteen tasks (three variations per task), and (2) a user study with nine participants to examine whether execution guards in the synthesized workflow can help users to understand and steer web agents. From the benchmark evaluation, web agents guided by the ReUseIt ’s synthesized workflows achieved a 70.1% success rate when executing repetitive tasks, improving the performance by 45.9% over executing without the workflows. From the user study, the execution guards effectively helped users to identify and understand agent issues, and boosted their willingness and trust to use web agents in their own tasks. In sum, this work makes three main contributions:
@@ 2. Related Work
LLM-based web agents and their challenges. LLM-based web agents are autonomous systems that leverage LLMs to navigate websites and execute tasks based on natural language instructions. These agents translate the LLM’s text outputs into web actions (e.g., click links and buttons, fill in forms, or web search) by specialized prompting and tool-use frameworks (OpenAI, 2021, 2025b). The usage of these agents has been explored in a wide range of scenarios, aimed at both professionals and end-users. On the professional side, web agents are being tested as digital “office workers” that could automate business workflows. For example, a recent benchmark, TheAgentCompany (Xu et al., 2025), simulated an enterprise intranet environment with internal web apps and data and evaluated LLM-based agents on everyday workplace tasks. On the consumer side, web agents can serve as advanced personal assistants that perform common web tasks on behalf of users. For instance, a web agent can book hotel rooms, search flight tickets, and reserve restaurants (Yang et al., 2024; Zhou et al., 2023). More use cases include tracking product prices on e-commerce websites, retrieving answers from multiple forums, and summarizing daily feeds on social media (Uzair, 2023; OpenAI, 2025a).
@@
Despite rapid progress, LLM-based web agents face significant challenges in repetitive tasks. A common concern is the reliability of task performance. Recent studies have found that only about 30% of multi-step web tasks can be completed successfully by advanced agents like Operator (OpenAI, 2025b), and many other agents achieve far lower success rates on average, often below 20% on realistic web automation benchmarks (Ning et al., 2025; Xue et al., ; Singh, 2025; Song et al., 2025). Another fundamental issue is inconsistent behavior that limits reusability. Since agent decisions emerge from probabilistic language model outputs, the exact sequence of actions can vary from run to run, and these agents fall short in being reused in similar tasks (Xue et al., ; OpenAI, 2025b; Pan et al., 2024; Levy et al., 2024). Such ad-hoc behavior also leads to unpredictable failures. When an unexpected webpage status change happens (e.g., a pop-up menu), the agent behaves differently across runs and fails in various ways (Ledel, 2025). To address these limitations, we focus on augmenting agent reliability and consistency with execution guards and self-recovery, and boosting reusability through successful and failed experiences from multiple executions of the same and similar tasks.
@@
Workflow synthesis for web agents. A promising approach to address web agents’ limitations is to automatically learn and reuse execution workflows as part of their reasoning process. For example, Agent Workflow Memory (AWM) (Wang et al., 2024b) induces commonly used workflows from an agent’s past experiences and injects them into the agent’s memory to guide future actions. Leveraging learned workflows, AWM significantly boosted success rates on long-horizon web tasks (e.g., +24.6% on the Mind2Web benchmark (Deng et al., 2023)) while reducing the number of execution steps. Other work has explored representing workflows in various forms to guide the agent, including natural language (Li et al., 2024), descriptive workflow language (Deng et al., 2023), and pseudo-code (Zhou et al., 2023; Wang et al., 2024b).
Nevertheless, the reusability of execution workflows has been less explored by prior research. Most of these studies analyze the task success rate of agents equipped with synthesized workflows, but offer limited insights into agent performance when reusing these workflows in repetitive and similar tasks. In our work, we induce agent workflows automatically, with a specific focus on synthesizing execution guards in the workflow to enhance agent reliability and reusability using any representation of the workflow.
@@
Programming by demonstration for web automation. Programming by demonstration (PbD) techniques have also been applied to address web automation challenges via scripting. In this paradigm, a human user performs a web task step-by-step, and the system synthesizes a deterministic program that replicates those actions. For example, WebRobot (Dong et al., 2022) constructs web automation scripts by observing user interactions with a web browser. DiLogics (Pu et al., 2023) leverages natural language processing to generalize user demonstrations into web automation scripts that handle diverse input conditions. MIWA (Chen et al., 2023) introduces a mixed-initiative interface for web automation, providing visual explanations and interactive debugging tools to help users understand and refine the generated scripts. In particular, programming by navigation synthesis (Lubin et al., 2025) has been proposed to support the iterative user specification refinement process.
@@
User interaction with AI agents. To enable users to inspect, interrupt, and redirect AI agents, researchers have explored interaction design that keeps humans in the loop during agent execution. Recent systems, such as Magentic-UI (Mozannar et al., 2025a), Cocoa (Feng et al., 2024), and Plan-Then-Execute (He et al., 2025), exemplify human interaction with web agents. They allow users to co-plan with web agents, observe real-time agent actions, take over tasks when agents get stuck, and approve/reject critical agent executions. In addition to web agents, recent research has improved the performance of and studied human interaction with agents in other types of graphical user interfaces (GUIs), such as mobile UIs (Zhang et al., 2025; Wang et al., 2024a; Hao et al., 2025; Yang et al., 2025), specific applications (Li et al., 2023; Liang et al., 2025), computer operating systems (Wu et al., 2024; Xie et al., 2024), and games (Wang et al., 2023). By investigating the impact of human involvement during AI agent execution, prior work has found improved user understanding, preference alignment, and agent safety. For instance, Uncertainty-Aware GUI Agent (Hao et al., 2025) can detect agent uncertainty in decision phases and request user feedback in ambiguous states; the Dark Patterns study (Tang et al., 2025b) shows that human oversight can help reduce errors when agents face deceptive interfaces; the Human-Centered Evaluation framework (Chen et al., 2025) suggests incorporating human evaluators into GUI agent assessment pipelines to ensure trustworthiness and alignment.
Following this idea, our work loops users into the agent’s execution process, such that users can identify and understand agent issues as they occur. Specifically, when the agent is uncertain about how to proceed or encounters a failure that it cannot recover from, it invites user guidance with issue analysis, aiming for more interpretable, trustworthy, and controllable agents.
@@ 3. Motivation
3.1. Preliminary Evaluation: Reliability of web agents to execute common web automation tasks
To quantify overall agent performance on common web automation tasks, we collected six tasks from prior work (He et al., 2024; Deng et al., 2025; Krosnick and Oney, 2021; Dong et al., 2022). These tasks require web agents to navigate the web, fill forms, apply filters, extract information, and summarize content, and the tasks span different application domains, including searching flights, housing, pets, online shopping products, research publications, and news. More details of the tasks are in Appendix A.1. We evaluated the web agent’s performance to complete these tasks under three conditions, to understand the agent’s ability to finish the tasks and reapply previous low-level action traces and high-level plans towards new tasks:
  • Condition 1 Task-Only: The agent is only provided with the task prompt from the user (e.g., “help me search for product information from an online shopping website”);
  • Condition 2 Task + Success-Traces: The agent is provided with the user prompt describing the task together with a previous successful execution trace containing low-level, detailed actions taken on each webpage (e.g., opened a drop-down menu, typed content in a text field; Figure 3 A);
  • Condition 3 Task + Magentic-UI Plan: The agent is provided with the task prompt together with a high-level workflow outlining the main steps obtained with Magentic-UI’s plan-learning method (e.g., step 1: navigate to a user-specified website, step 2: search content; Figure 3 B).
This figure shows three examples of action traces/workflows. On the left, it shows ”A Success-Traces”, where a list of detailed agent actions is outlined, such as text typed and buttons clicked. In the middle, it displays ”B Magentic-UI Plan”, where a list of high-level steps containing the main actions to take is shown, such as navigating to a website and entering user-specified information. On the right, it illustrates ”C Magentic-UI Plan User”, which is largely the same as ”B Magentic-UI Plan” but with user edits, such as action cues to help the agent recover from failures.
To measure the agent’s performance, we used the success rate ($SR$) as the evaluation metric. $SR$ is defined as the number of successful executions out of the total number of runs. For each task, we ran five executions and computed the fraction of successful runs. We also generated variation tasks (with variations on task attribute, category, and website; more details are presented in Section 4.1) and executed each variation five times to test workflow reuse on closely related tasks. In total, we executed each task and its variations 20 times, and report the mean $\pm$ std $SR$ (in $\%$) across all these executions. To determine whether the executions were successful, we manually inspected the screen recordings to judge if the agent had satisfied all user requirements and found the intended results.